OpenStudy (anonymous):

y + 5 = –12 (1 point) y = –19 y = –17 y = –7 y = 7

jimthompson5910 (jim_thompson5910):

Subtract 5 from both sides

OpenStudy (anonymous):

so what is -12 - 5

OpenStudy (amorfide):

y + 5 = –12 you want to get all constants onto one side and all unknowns onto another side to do this you look at the terms and decide which is easier, well in this case we should just move the 5 from the left to the right, to do this you do the opposite of the sign infront of the constant, so we have +5, you want to subtract 5 but anything done to the left must be done to the right so you rewrite this to be y+5-5=-12-5 y=-17